# Probing the accretion geometry of black hole X-ray binaries: A multi-mission spectro-polarimetric and timing study

### v1

@@ 5.2 Model Prescriptions and Limitations of X-ray Polarization in BH-XRBs
X-ray polarization measurements with IXPE reveal significant polarized emission across several spectral states of the sources in $2-8$ keV energy range, except for LMC X--1 (see §4.3 and Podgorný et al. 2023), Swift J151857.057214151857.057214151857.0-57214151857.0 - 57214 (see §4.3 and Ling et al. 2024) and MAXI J174429417442941744-2941744 - 294 (see §4.3). We find that Cyg X--1 exhibits PD within $3-4.8\%$ in LHS, which decreases to $1.4-2.5\%$ in HSS. A similar noticeable reduction in the PD from $4.7\%$ (HIMS) to $0.4\%$ (HSS) is also observed for the exceptionally bright transient Swift J1727.8--1613. The most recent IXPE observation of this source confirms the recovery of PD $\sim 3.2\%$ in the faint LHS. We mention that these results are in agreement with the previously reported polarimetric findings of Cyg X--1 (Krawczynski et al., 2022; Jana & Chang, 2024; Steiner et al., 2024) and Swift J1727.8--1613 (Veledina et al., 2023; Ingram et al., 2024; Svoboda et al., 2024b; Podgorný et al., 2024).
@@
Intriguingly, an exceptionally high PD of $\sim 8.3\%$ is observed in the HSS of 4U 16304747-47- 47, which drops to $\sim 6.8\%$ during the SPL state exhibiting strong ($>4\sigma$) energy dependency (see Table 5 and Fig. 5; also Kushwaha et al. 2023b; Rodriguez Cavero et al. 2023; Rawat et al. 2023b, a; Ratheesh et al. 2024). In contrast, low PD values are observed in several sources: LMC X33-3- 3 ($\sim 2.4-3\%$), 4U 1957$+115$ ($\sim 1.9\%$) and GX $339-4$ ($\sim 1.2\%$), being consistent with the previous measurements (Majumder et al., 2024a; Svoboda et al., 2024a; Garg et al., 2024; Marra et al., 2024; Mastroserio et al., 2025). In the present work, we report polarization detections analyzing the most recent IXPE observations of Cyg X--3 (PD $\sim 21.4\%$, SIMS), LMC X--3 (PD $\sim 2.4\%$, HSS) and IGR J17091--3624 (PD $\sim 9\%$, LHS), apart from the earlier measurements of these sources. In addition, for the first time to the best of our knowledge, we report the null-detection of X-ray polarization in LMC X33-3- 3 (E3) and a newly discovered BH-XRB candidate MAXI J174429417442941744-2941744 - 294.
@@
Meanwhile, several models have been proposed to account for the observed polarization in BH-XRBs. For instance, remarkably high PD ($\sim 8.3\%$) of 4U 1630--47 is attributed to the emission from a partially ionized thin/slim accretion disc viewed at inclination $i\gtrsim 60^{\circ}$ (Ratheesh et al., 2024). Similar explanation is also suggested for LMC X33-3- 3 in the HSS that exhibits PD $\sim 3\%$ at $i\sim 70^{\circ}$ (Majumder et al., 2024a; Svoboda et al., 2024a). Further, scattering of photons in strong disc winds of different opening angles ($\alpha_{\rm w}\sim 30^{\circ}-40^{\circ}$) has also been suggested as another possible mechanism to explain high PD values of 4U 1630--47 (Kushwaha et al., 2023b; Rawat et al., 2023b; Nitindala et al., 2025). On the other hand, in high-spin black hole systems, strong relativistic effects cause intense gravitational lensing that bends emitted photons back onto the disc. This enhances the reflection features and significantly affects the observed polarization (see §4.2, Schnittman & Krolik, 2009). Such an effect appears consistent with the polarimetric observations (PD $\sim 1.2-2.5\%$) in the soft state as reported for rapidly spinning systems like Cyg X11-1- 1 ($a_{*}>0.998$) and 4U $1957+115$ ($a_{*}\sim 0.988$) (Marra et al., 2024; Steiner et al., 2024).
@@
Furthermore, it has been suggested that $\sim 4\%$ (Cyg X11-1- 1, LHS) and $\sim 9\%$ (IGR J17091362417091362417091-362417091 - 3624, LHS) polarization can be obtained from a wedge-shaped corona with mildly relativistic outflowing plasma having inclination in the range $30^{\circ}\lesssim i\lesssim 60^{\circ}$ (Poutanen et al., 2023; Ewing et al., 2025). A similar model of bulk Comptonization involving a mildly relativistic outflow predicts an even higher polarization degree up to $\lesssim 10\%$ (Dexter & Begelman, 2024). Notably, observed $\sim 4.7\%$ PD at $i\sim 40^{\circ}$ in HIMS of Swift J1727.816131727.816131727.8-16131727.8 - 1613 also seems to be consistent with the predictions from a wedge-shaped corona with outflowing plasma velocity $\lesssim 0.2c$ (Poutanen et al., 2023). However, QPOs are detected in Swift J1727.816131727.816131727.8-16131727.8 - 1613 ($0.5-100$ keV) and IGR J17091362417091362417091-362417091 - 3624 ($3-79$ keV) during observations simultaneous with IXPE, although their physical origin remains unclear within the framework of the outflowing corona geometry. On the other hand, Kumar (2024) predicted PD $\sim 0.5-4\%$ assuming a simple static spherical corona for $30^{\circ}\lesssim i\lesssim 60^{\circ}$. This appears broadly consistent with the low PD of $\sim 1.2\%$ observed at $i\sim 30^{\circ}$ in the SIMS of GX $339-4$.
@@ 5.3 Accretion-Ejection Scenarios in BH-XRBs
Figure 7: Schematic representation of the possible disc-corona-jet configurations for seven BH-XRB sources based on the framework of two-component accretion flow scenario. The Keplerian and sub-Keplerian components, as indicated in panel (b), are also present in the other two configurations. Colored vertical arrows in the corona regions of panels (a) and (c) indicate the direction of polarization angle for the respective sources. In panels (a) and (c), the jet configuration applies only to the sources in distinct spectral states, which are marked by their respective serial numbers. See the text for details.
@@
Following the discussion of model degeneracies and the need for a unified model framework, we next explore physically motivated accretion-ejection geometries capable of explaining the commonly observed spectro-temporal features, while also offering a promising basis for interpreting the spectro-polarimetric signatures in BH-XRBs. Towards this, a plausible disc-corona-jet configuration based on the two-component accretion flow model Chakrabarti & Titarchuk (1995) is illustrated in Fig. 7 and discussed in detail below.
@@
The spectro-polarimetric correlation study (see §4.6 and Fig. 6) suggests that the polarization signatures (PD $\sim 3-9\%$) in the LHS/HIMS of Cyg X11-1- 1, Swift J1727.816131727.816131727.8-16131727.8 - 1613 and IGR J17091362417091362417091-362417091 - 3624 are primarily governed by the dominant Comptonization process characterized by $F_{\rm ratio}<1$ and $cov_{\rm frac}\sim 35-87\%$. Interestingly, the radio jet position angle of Cyg X11-1- 1 and Swift J1727.816131727.816131727.8-16131727.8 - 1613 roughly aligns along the observed PA throughout the respective IXPE campaigns (Miller-Jones et al., 2021; Wood et al., 2024). This suggests that PA is possibly oriented perpendicular to the disc-corona geometry. Further, it may be noted that PA resulting from multiple Compton up-scatterings generally tends to align along the minor axis of the corona (Ingram et al., 2024). These pieces of evidence suggest a disc-corona configuration comprising a radial corona located close to the black hole, along with a truncated accretion disc at larger radius during LHS/HIMS (see Fig. 7a) of Swift J1727.816131727.816131727.8-16131727.8 - 1613 and Cyg X11-1- 1. The above configuration potentially explains the QPOs in Swift J1727.816131727.816131727.8-16131727.8 - 1613 (see §4.2 and Fig. 2c) in terms of the oscillating boundary of the Comptonizing region (see Fig. 7a and Molteni et al. 1996; Nandi et al. 2024). It is worth mentioning that the coronal geometry in IGR J17091362417091362417091-362417091 - 3624 remains inconclusive from the polarimetric results due to the unresolved position angle of the radio jet. Although the presence of QPO in the LHS of IGR J17091362417091362417091-362417091 - 3624 (see Fig. 2g) perhaps indicates a similar disc-corona configuration presented in Fig. 7a. However, the absence of QPO in Cyg X11-1- 1 implies that either the coronal boundary fails to satisfy the necessary resonance condition for modulation, or the oscillation is too subtle to be detected as predicted by Majumder et al. (2025a).
@@
For 4U $1630-47$, an absorption feature around $\sim 7$ keV is observed, suggesting the presence of strong disc winds in the HSS ($F_{\rm ratio}\sim 52$ and $cov_{\rm frac}\sim 5\%$) spectra (see §4.2). Accordingly, a disc-dominated accretion configuration along with a weak corona, accompanied by strong disc winds with varying opening angles (see also Nitindala et al. 2025), is likely favored for this source (see Fig. 7b). However, a relatively strong corona ($cov_{\rm frac}\sim 38\%$) may be present in the SPL state as compared to the HSS (see Fig. 7b).
@@
For GX $339-4$, a low PD ($\sim 1.2\%$) is observed within $3-8$ keV energy range during SIMS ($cov_{\rm frac}\sim 34\%$ and $F_{\rm ratio}\sim 2.4$), while null-detection of PD is seen in the HSS (Table 5). This possibly associates the polarization signature of the source with Comptonized emission. We notice $\rm PD<\rm MDP_{99}\%$ in the entire IXPE energy range of $2-8$ keV during SIMS, which possibly resulted from the depolarization of the overall radiation by unpolarized and/or oppositely polarized (as compared to corona) disc emission (Ingram et al., 2024) at lower energies ($2-3$ keV). The polarization angle of GX $339-4$ aligns with the direction of the discrete radio jet knot (Mastroserio et al., 2025). These findings predict the GX $339-4$ configuration comprising a radially extended weak compact corona coexisting with a strong disc ($F_{\rm ratio}\sim 2.4$, $\Gamma_{\rm th}\sim 2$), as illustrated in Fig. 7c. Moreover, a type-B QPO (see Fig. 2e) observed in this source could be explained from this configuration using an analogous argument presented above for Fig. 7a (see also Nandi et al. 2012; Aneesha et al. 2024).
@@
Furthermore, we conjecture that a similar disc-corona configuration (see Fig. 7c) is also preferred in the disc-dominated HSS/SIMS of other sources (LMC X33-3- 3, 4U $1957+115$, Cyg X11-1- 1 and Swift J1727.816131727.816131727.8-16131727.8 - 1613) with major contributions from thermal disc and/or various general relativistic effects (Majumder et al., 2024a; Svoboda et al., 2024a; Marra et al., 2024; Steiner et al., 2024; Svoboda et al., 2024b) in the observed PD ($1.4-3\%$). It is noteworthy that although the presence of a radial corona is favored for 4U $1630-47$, LMC X33-3- 3, and 4U $1957+115$ in Fig. 7, the coronal geometry remains uncertain due to the lack of resolved radio jet orientation for these sources. Moreover, high PD (up to $\sim 21\%$) dominated by reflected emission in Cyg X33-3- 3 appears inconsistent with the disc-corona configurations presented in Fig. 7. It has been suggested that a disc-corona geometry featuring an optically thick funnel in the innermost region of the accretion disc, which obscures most of the primary emission, is favored for Cyg X33-3- 3 Veledina et al. (2024a, b). Finally, we mention that, although the two-component accretion flow configuration aligns with the disc-corona geometry inferred from polarimetric studies for most of the BH-XRBs, a detailed radiative transfer computation is needed within this framework for the predictions on X-ray polarization.

### v2

@@ 5.2 Model Prescriptions and Limitations of X-ray Polarization in BH-XRBs
X-ray polarization measurements with IXPE reveal significant polarized emission across several spectral states of the sources in 282-8 keV energy range, except for LMC X-1 (see §4.3 and Podgorný et al. 2023), Swift J151857.057214151857.0-57214 (see §4.3 and Ling et al. 2024) and MAXI J17442941744-294 (see §4.3). We find that Cyg X-1 exhibits PD within $3-4.8\%$ in LHS, which decreases to $1.4-2.5\%$ in HSS. A similar noticeable reduction in the PD from $4.7\%$ (HIMS) to $0.4\%$ (HSS) is also observed for the exceptionally bright transient Swift J1727.8-1613. The most recent IXPE observation of this source confirms the recovery of PD $\sim 3.2\%$ in the faint LHS. We mention that these results are in agreement with the previously reported polarimetric findings of Cyg X-1 (Krawczynski et al., 2022; Jana & Chang, 2024; Steiner et al., 2024) and Swift J1727.8-1613 (Veledina et al., 2023; Ingram et al., 2024; Svoboda et al., 2024b; Podgorný et al., 2024).
@@
Intriguingly, an exceptionally high PD of $\sim 8.3\%$ is observed in the HSS of 4U 163047-47, which drops to $\sim 6.8\%$ during the SPL state exhibiting strong ($4\sigma$) energy dependency (see Table 5 and Fig. 5; also Kushwaha et al. 2023; Rodriguez Cavero et al. 2023; Rawat et al. 2023b, a; Ratheesh et al. 2024). In contrast, low PD values are observed in several sources: LMC X3-3 ($\sim 2.4-3\%$), 4U 1957+115+115 ($\sim 1.9\%$) and GX 3394339-4 ($\sim 1.2\%$), being consistent with the previous measurements (Majumder et al., 2024a; Svoboda et al., 2024a; Garg et al., 2024; Marra et al., 2024; Mastroserio et al., 2025). In the present work, we report polarization detections analyzing the most recent IXPE observations of Cyg X-3 (PD $\sim 21.4\%$, SIMS), LMC X-3 (PD $\sim 2.4\%$, HSS) and IGR J17091-3624 (PD $\sim 9\%$, LHS), apart from the earlier measurements of these sources. In addition, for the first time to the best of our knowledge, we report the null-detection of X-ray polarization in LMC X3-3 (E3) and a newly discovered BH-XRB candidate MAXI J17442941744-294. We note that, subsequent to our report, null-detection of polarization for MAXI J17442941744-294 is also confirmed by Marra et al. (2025).
@@
Meanwhile, several models have been proposed to account for the observed polarization in BH-XRBs. For instance, remarkably high PD ($\sim 8.3\%$) of 4U 1630-47 is attributed to the emission from a partially ionized thin/slim accretion disc viewed at inclination $i\gtrsim 60^{\circ}$ (Ratheesh et al., 2024). Similar explanation is also suggested for LMC X3-3 in the HSS that exhibits PD $\sim 3\%$ at $i\sim 70^{\circ}$ (Majumder et al., 2024a; Svoboda et al., 2024a). Further, scattering of photons in strong disc winds of different opening angles ($\alpha_{\rm w}\sim 30^{\circ}-40^{\circ}$) has also been suggested as another possible mechanism to explain high PD values of 4U 1630-47 (Kushwaha et al., 2023; Rawat et al., 2023b; Nitindala et al., 2025). However, Tomaru et al. (2024) demonstrated that electron scattering in an optically thin wind produces polarization in a direction opposite to that obtained from the intrinsic emission from an optically thick disc. As a result, the wind component is expected to depolarize the total emission, including the intrinsic disc emission. Therefore, the presence of disc-wind from the optically thick disc appears to be insufficient to account for the observed high PD in 4U 1630-47. On the other hand, in high-spin black hole systems, strong relativistic effects cause intense gravitational lensing that bends emitted photons back onto the disc. This enhances the reflection features and significantly affects the observed polarization (see §4.2, Schnittman & Krolik, 2009). Such an effect appears consistent with the polarimetric observations (PD $\sim 1.2-2.5\%$) in the soft state as reported for rapidly spinning systems like Cyg X1-1 ($a_{*}>0.998$) and 4U $1957+115$ ($a_{*}\sim 0.988$) (Marra et al., 2024; Steiner et al., 2024).
@@
Furthermore, it has been suggested that $\sim 4\%$ (Cyg X1-1, LHS) and $\sim 9\%$ (IGR J17091362417091-3624, LHS) polarization can be obtained from a wedge-shaped corona with mildly relativistic outflowing plasma model of Beloborodov (1998) having inclination in the range $30^{\circ}\lesssim i\lesssim 60^{\circ}$ (Poutanen et al., 2023; Ewing et al., 2025). A similar model of bulk Comptonization involving a mildly relativistic outflow predicts an even higher polarization degree up to $\lesssim 10\%$ (Dexter & Begelman, 2024). Notably, observed $\sim 4.7\%$ PD at $i\sim 40^{\circ}$ in HIMS of Swift J1727.816131727.8-1613 also seems to be consistent with the predictions from a wedge-shaped corona with outflowing plasma velocity $\lesssim 0.2c$ (Poutanen et al., 2023). However, QPOs are detected in Swift J1727.816131727.8-1613 (0.51000.5-100 keV) and IGR J17091362417091-3624 (3793-79 keV) during observations simultaneous with IXPE, although their physical origin remains unclear within the framework of the outflowing corona geometry. On the other hand, Kumar (2024) predicted PD $\sim 0.5-4\%$ assuming a simple static spherical corona for $30^{\circ}\lesssim i\lesssim 60^{\circ}$. This appears broadly consistent with the low PD of $\sim 1.2\%$ observed at $i\sim 30^{\circ}$ in the SIMS of GX 3394339-4.
@@ 5.3 Accretion-Ejection Scenarios in BH-XRBs
Figure 7: Schematic representation of the possible disc–corona–jet configurations for seven BH-XRB sources within the two-component accretion flow framework. The Keplerian and sub-Keplerian components shown in panel (b) are also present in the other three configurations. Colored vertical arrows in the corona regions of panels (a) and (c) denote the polarization angle for the respective sources. In panels (a) and (d), the jet/radio emission configuration applies only to sources in distinct spectral states, indicated by their serial numbers. See the text for details.
@@
Following the discussion of model degeneracies and the need for a unified model framework, we next explore physically motivated accretion-ejection geometries capable of explaining the commonly observed spectro-temporal features, while also offering a promising basis for interpreting the spectro-polarimetric signatures in BH-XRBs. Towards this, a plausible disc-corona-jet configuration based on the two-component accretion flow model Chakrabarti & Titarchuk (1995); Chakrabarti (1996) is illustrated in Fig. 7. In this model, the accretion flow around black holes comprise with Keplerian disc flanked between the sub-Keplerian components. This sub-Keplerian component generally forms the post-shock corona (PSC, equivalently Comptonizing corona) in the inner region of the disc Chakrabarti & Titarchuk (1995); Molteni et al. (1996); Chakrabarti et al. (2008); Nandi et al. (2012); Das et al. (2014). Soft photons from the Keplerian disc are up-scattered by hot electrons present in the corona and produce the hard spectral tail seen in BH-XRBs. The inner edge of the Keplerian disc smoothly connects to the outer boundary of the coronal region.
@@
The spectro-polarimetric correlation study (see §4.6 and Fig. 6) suggests that the polarization signatures (PD $\sim 3-9\%$) in the LHS/HIMS of Cyg X1-1, Swift J1727.816131727.8-1613 and IGR J17091362417091-3624 are primarily governed by the dominant Comptonization process characterized by $F_{\rm ratio}<1$ and $cov_{\rm frac}\sim 35-87\%$. Interestingly, the radio jet position angle of Cyg X1-1 and Swift J1727.816131727.8-1613 roughly aligns along the observed PA throughout the respective IXPE campaigns (Miller-Jones et al., 2021; Wood et al., 2024). This suggests that PA is possibly oriented perpendicular to the disc-corona geometry. Further, it may be noted that PA resulting from multiple Compton up-scatterings generally tends to align along the minor axis of the corona (Ingram et al., 2024). These pieces of evidence suggest a disc-corona configuration comprising a radial corona located close to the black hole, along with a truncated accretion disc at larger radius during LHS/HIMS (see Fig. 7a) of Swift J1727.816131727.8-1613 and Cyg X1-1. The above configuration potentially explains the QPOs in Swift J1727.816131727.8-1613 (see §4.2 and Fig. 2c) in terms of the oscillating boundary of the Comptonizing region (see Fig. 7a and Molteni et al. 1996; Nandi et al. 2024). It is worth mentioning that the coronal geometry in IGR J17091362417091-3624 remains inconclusive from the polarimetric results due to the unresolved position angle of the radio jet. Although the presence of QPO in LHS of IGR J17091362417091-3624 (see Fig. 2g) perhaps indicates a similar disc-corona configuration presented in Fig. 7a. However, the absence of QPO in Cyg X1-1 implies that either the coronal boundary fails to satisfy the necessary resonance condition for modulation, or oscillation is too subtle to be detected as predicted by Majumder et al. (2025a).
@@
For 4U 1630471630-47, an absorption feature around $\sim 7$ keV is observed, suggesting the presence of strong disc winds in the HSS ($F_{\rm ratio}\sim 52$ and $cov_{\rm frac}\sim 5\%$) spectra (see §4.2). Accordingly, a disc-dominated accretion configuration along with a weak corona, accompanied by strong disc winds with varying opening angles (see also Nitindala et al. 2025), is likely favored for this source (see Fig. 7b). However, a relatively strong corona ($cov_{\rm frac}\sim 38\%$) may be present in the SPL state as compared to the HSS (see Fig. 7b).
@@
For GX 3394339-4, a low PD ($\sim 1.2\%$) is observed within 383-8 keV energy range during SIMS ($cov_{\rm frac}\sim 34\%$ and $F_{\rm ratio}\sim 2.4$), while null-detection of PD is seen in the HSS (Table 5). This possibly associates the polarization signature of the source with Comptonized emission. We notice $\rm PD<\rm MDP_{99}\%$ in the entire IXPE energy range of 282-8 keV during SIMS, which possibly resulted from the depolarization of the overall radiation by unpolarized and/or oppositely polarized (as compared to corona) disc emission (Ingram et al., 2024) at lower energies (232-3 keV). The polarization angle of GX 3394339-4 aligns with the direction of the discrete radio jet knot (Mastroserio et al., 2025). These findings predict the GX 3394339-4 configuration comprising a radially extended weak compact corona coexisting with a strong disc ($F_{\rm ratio}\sim 2.4$, $\Gamma_{\rm th}\sim 2$), as illustrated in Fig. 7c. Moreover, a type-B QPO (see Fig. 2e) observed in this source could be explained from this configuration using an analogous argument presented above for Fig. 7a (see also Nandi et al. 2012; Aneesha et al. 2024). Notably, the SIMS of Swift J1727.816131727.8-1613 also corresponds to the configuration shown in Fig. 7c.
@@
Furthermore, we conjecture that a similar disc-corona configuration (see Fig. 7d) is also preferred in the disc-dominated HSS of other sources (LMC X3-3, 4U $1957+115$, Cyg X1-1 and Swift J1727.816131727.8-1613) with major contributions from thermal disc and/or various general relativistic effects (Majumder et al., 2024a; Svoboda et al., 2024a; Marra et al., 2024; Steiner et al., 2024; Svoboda et al., 2024b) in the observed PD ($1.4-3\%$). Among these, quenched radio emission with flux densities of $\sim 2.4-11.7$ mJy is observed in the HSS of Cyg X1-1 (see Table 1 and Fig. 7d), which is found to be consistent with previous reports of radio quenching in this source (Tigelaar et al., 2004). It is noteworthy that although the presence of a radial corona is favored for 4U 1630471630-47, LMC X3-3, and 4U $1957+115$ in Fig. 7, the coronal geometry remains uncertain due to the lack of resolved radio jet orientation for these sources. Moreover, high PD (up to $\sim 21\%$) dominated by reflected emission in Cyg X3-3 appears inconsistent with the disc-corona configurations presented in Fig. 7. It has been suggested that a disc-corona geometry featuring an optically thick funnel in the innermost region of the accretion disc, which obscures most of the primary emission, is favored for Cyg X3-3 Veledina et al. (2024a, b). Finally, we mention that, although the two-component accretion flow configuration aligns with the disc-corona geometry inferred from polarimetric studies for most of the BH-XRBs, a detailed radiative transfer computation is needed within this framework for the predictions on X-ray polarization. Therefore, the schematic in Fig. 7 illustrates a possible disc-corona-jet geometry in BH-XRBs; however, it does not yet provide a quantitative explanation for the polarimetric results.